an object thrown up with a instant velocity 40m/s from grounds. assume the energy is conservative. find

a) maximum height

b)when where does Porential energy= Kinetic Energy if mass m=2kg


1
Expert's answer
2021-01-11T11:35:24-0500

"E_p =mgh\text { where }g=9.8m/s^2"

"E_k=\frac{mV^2}{2}"

"\text{if }h =0 \text{ then }V_0=40\text{ by problem statement}:"

"E_p+E_k= mg*0+\frac{mV_0^2}{2}=\frac{mV_0^2}{2}"

"\frac{mV_0^2}{2} - \text{this value is constant for } E_p+E_k"

"a) \text{for }h_{max}\ V=0:"

"E_p+E_k= mgh_{max}+\frac{m*0}{2}= mgh_{max}"

"mgh_{max}=\frac{mV_0^2}{2}"

"h_{max}=\frac{V_0^2}{2g}=\frac{40^2}{2*9.8}\approx81.6"

"b)E_p=E_k"

"2E_p = \frac{mV_0^2}{2}"

"E_p = \frac{mV_0^2}{4}= \frac{2*40^2}{4}=800"

"E_k=800"

Answer: "a)h_{max}\approx81.6m\ b)\text{if }E_p=E_k \text{ then}E_p=E_k=800J"